OpenStudy (anonymous):

Is the simplified form of 2square root of 3 • square root of 12 rational? Yes No

OpenStudy (anonymous):

help

OpenStudy (solomonzelman):

\[2\sqrt{3} \times \sqrt{12}\]\[2\sqrt{3}\times(\sqrt{4}\times \sqrt{3})\]\[2\sqrt{3} \times (2\sqrt{3})\]\[4\times 3\] rational or irrational?

OpenStudy (anonymous):

rational right

OpenStudy (solomonzelman):

Yup!

OpenStudy (solomonzelman):

All you need for this type of questions is to see how to brake the square roots, see how I broke sqrt12? that's the key....

OpenStudy (anonymous):

oh okay so its yes

OpenStudy (solomonzelman):

Yes, it's yes.

OpenStudy (anonymous):

ok lol thanks

OpenStudy (anonymous):

Simplify: square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses 14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above

OpenStudy (solomonzelman):

Please post one question per thread, I will help you with this, but next time please one question per thread.

OpenStudy (solomonzelman):

Write it out for me.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt{5}(10-4\sqrt{2})\]

OpenStudy (solomonzelman):

lets expand he parenthesis, \[10\sqrt{5}-4\sqrt{10}\] and this is the final answer.

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

wich onw was it A B C or D

OpenStudy (solomonzelman):

Well it's not A, B, or C.

OpenStudy (anonymous):

SO ITS D !!!!